OpenStudy (anonymous):

ER = (2a + b) inches. What is the area of square OSER?

OpenStudy (anonymous):

 this is the figure

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

Area of Square A = s^2 A = (2a+b)^2 A = ???

OpenStudy (anonymous):

These are the answer choices. A. (4a squared + 2ab + b squared) inches squared B. (4a squared + 2b squared) inches squared C. (4a squared + b squared) inches squared D. (4a + 2b) inches squared E. (4a squared + 4ab + b squared) inches squared

jimthompson5910 (jim_thompson5910):

(2a+b)^2 = (2a+b)(2a+b) = ???

OpenStudy (anonymous):

I need help please @aum @Bram @Broskishelleh @BVB4life @IndiiAvery @LifeEngineer @LearningIsAwesome @lxl-rhea-lxl

jimthompson5910 (jim_thompson5910):

Are you able to expand out (2a+b)(2a+b) ?

OpenStudy (anonymous):

I tried with no avail .-. i'm sorry :(

OpenStudy (anonymous):

@aum can you help me out??